The following information is available regarding the total manufacturing overhead of Bursa Mfg. Co. for a recent four-month period.

Machine-
Hours
Manufacturing Overhead
January 5,100 $ 300,000
February 3,200 224,000
March 4,900 263,800
April 2,900 180,000


a-1. Use the high-low method to determine the variable element of manufacturing overhead costs per machine-hour. (Round your answer to 2 decimal places.)

a-2. Use the high-low method to determine the fixed element of monthly overhead cost.

b. Bursa expects machine-hours in May to equal 5,300. Use the cost relationships determined in part a to forecast May's manufacturing overhead costs.

c. Suppose Bursa had used the cost relationships determined in part a to estimate the total manufacturing overhead expected for the months of February and March. By what amounts would Bursa have over- or underestimated these costs?
Answer #1

a - 1

Observation of given data indicates that, Month of January is high volume month and April is low.

M.Hours Overhead
Jan 5100 300000
April 2900 180000
Difference 2200 120000

Variable overhead per machine hour = Difference in machine hours / Difference in Machine hours

= 120,000 / 2200

Final answer =  54.55

a - 2

Fixed manufacturing Overhead = Total overhead - Variable overhead

= 300,000 - 5100 * 54.55

= 21795

Question - (b)

Forecast for the month of may = 21795 + 5300 * 54.55 = 310,910

Question - (c)

Forecast for the month of Feb = 21795 + 3200 * 54.55 = 196355

But actual overhead for Feb = 224000

Thus we have under-estimation of overhead for Feb = 224000 - 196355 = 27,645 Under estimation

Forecast for the month of March = 21795 + 4900 * 54.55 = 289090

But actual overhead for March = 263800

Thus we have Over-estimation of overhead for Feb = 263800 - 289090 = 25290 Over estimation

(Note : If negative numbers are allowed ............ enter the over-estimation as - 25290 )